OpenStudy (anonymous):

Can anyone help me learn how to solve \[\sqrt{75}\]

OpenStudy (kc_kennylau):

Prime-factorise 75

OpenStudy (anonymous):

my bad i forgot some of the text, i need to learn how to simplify it.

OpenStudy (kc_kennylau):

Prime-factorize 75 first

OpenStudy (kc_kennylau):

That means break it down into primes

OpenStudy (anonymous):

ok... 1, 3, 5, 15 and 25

OpenStudy (anonymous):

my bad its 3x5

OpenStudy (kc_kennylau):

My bad I mean product of primes

OpenStudy (anonymous):

so 3x5 is the product of primes or is it 3 x 5 = 15 x 5 = 75

OpenStudy (kc_kennylau):

Okay it should be 3x5x5

OpenStudy (anonymous):

ok

OpenStudy (kc_kennylau):

\[\sqrt{75}=\sqrt{3\times5\times5}=\sqrt3\sqrt{5\times5}=5\sqrt3\]

OpenStudy (anonymous):

thank you, 1 more?

OpenStudy (kc_kennylau):

okay

OpenStudy (anonymous):

thnxs, Simply \[\sqrt{864}\]

OpenStudy (kc_kennylau):

Just follow what I did

OpenStudy (anonymous):

so it would be \[\sqrt{9(96)} = \sqrt{9(3)(32)} = \sqrt{9(3)(16)(2)} = 12\sqrt{6}?\]

OpenStudy (kc_kennylau):

Yes exactly

OpenStudy (anonymous):

YES! Thank you SOO much, you just saved me hours of homework! :D

OpenStudy (kc_kennylau):

No problem :)

OpenStudy (kc_kennylau):

I'm flattered :)